OpenStudy (anonymous):

Consider the differential equation 2y′′+2y′+2y=0 with y(0)=−1 , y′(0)=−3. Taking the Laplace transform and solving for L(y) yields : L(y)=F(s) where F(s)=

OpenStudy (anonymous):

I get F(s) = \[\frac{ -2s-8 }{ 2s ^{2}+2s + 2 }\] for an answer but it is wrong

zepdrix (zepdrix):

Hmmmm it looks correct... Everything has a 2 in it though, do they want it simplified maybe?

OpenStudy (usukidoll):

split the fraction up and simplify ... before inverse laplace?

OpenStudy (unklerhaukus):

\[2y''+2y'+2y=0,\qquad y(0)=-1,\quad y'(0)=3\\~\\ \quad\color{gray}{2\Big(}y''+y'+y\color{gray}{\Big)}=0\\ ~\\ \left[s^2Y-sy(0)-y'(0)\right]+\Big[sY-y(0)\Big]+\Big[Y\Big]=0\\ Y(s^2+s+1)-y(0)(s+1)-y'(0)=0\\ Y(s^2+s+1)+(s+1)+3=0\\ Y(s^2+s+1)=-(s+4)\\ Y=-\frac{s+4}{s^2+s+1}=-\frac{\color{gray}{2\big(}s+4\color{gray}{\big)}}{\color{gray}{2\big(}s^2+s+1\color{gray}{\big)}}\\ \]

OpenStudy (usukidoll):

oh I see where this is going... cancel the 2's out

OpenStudy (usukidoll):

and then do partial fractions b^2-4ac will tell you whether or not s^2+s+1 can be factorable

OpenStudy (usukidoll):

 ok so it's a complex root...it's best to leave the denominator alone or you'll be in a mess